Suppose w is a cube root of unity with w not equal to 1 suppose P and Q are the points on complex plane defined by w and (w^2) if O is the origin then what is angle between OP and OQ
Sir solve this problem from easy mathod

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

Suppose w is a cube root of unity with w not equal to 1 suppose P and Q are the points on complex plane defined by w and (w^2) if O is the origin then what is angle between OP and OQ
Sir solve this problem from easy mathod
----------
120 degrees
--------------
I think my method (knowing the answer) is easier than the other tutor's.

Answer by greenestamps(13200)



The n-th roots of unity are distributed evenly, with the angle in degrees between them being 360/n.

The angle between any two "successive" 3rd roots of unity is 360/3 = 120 degrees.
The angle between any two "successive" 4th roots of unity is 360/4 = 90 degrees. (The four 4th roots of unity are i, -1, -i, and 1).
The angle between any two "successive" 5th roots of unity is 360/5 = 72 degrees.
The angle between any two "successive" 6th roots of unity is 360/6 = 60 degrees.
etc., etc.,...